Silver tarnishes in the presence of hydrogen sulfide (which smells like rotten eggs) and oxygen because of the following unbalanced chemical reaction


Ag + H2S + O2 à Ag2S + H2O


How many grams of silver sulfide can be formed from 1.1 g Ag reacting with 1.7 g of H2S and 1.9 g of O2

1
Expert's answer
2021-06-24T13:08:03-0400

4Ag + 2H2S + O2 = 2Ag2S + 2H2O


"n(Ag)=\frac{m}{M}=frac{1.1}{107.8682}=0.0102"


"\frac{0.0102}{4}=\frac{x}{2}"


x=0.005


"m(Ag2S)=n\times M=0.005\times247.8024=1.26g"